OpenStudy (anonymous):

An industrial production process costs C(q) million dollars to produce q million units; these units then sell for R(q) million dollars. If C(2.1)=5.1 , R(2.1)=6.7 , MC(2.1)=0.4 and MR(2.1)=0.5, calculate the following. (a) The profit earned by producing 2.1 million units. ((I actually got this right already. 6.7-5.1 = 1.6)) (b) The approximate change in revenue if production increases from 2.1 to 2.14 million units. (c) The approximate change in revenue if production decreases from 2.1 to 2.07 million units. (d) The approximate change in profit in parts (b) and (c).

OpenStudy (anonymous):

So I need help on b, c, d.

OpenStudy (anonymous):

(b) Δq=0.03 ΔR (2.1) Δq = 0.5(0.03)=0.015= $15 thousand (c) Δq=-0.07 ΔR = R(2.1)Δq = 0.5(-0.07) = -0.035 = -$35 thousand (d) Δq=-0.07 ΔC = C(2.1)Δq = 0.4(-0.07) = -0.028 = $28,000 Δπ = -35000 - (-28000) = -$7000 This is what the computer gave me because I used up my 5 tries. I get the actual calculations, but where does Δq=0.03 for (b) and Δq=-0.07 for (c) and (d) come from?